# Rigorous Modeling of Droplet Impact Dynamics on Micro-Structured Surfaces: Nonlocal Theory and SPH Simulation of Pancake Bouncing

### v1

@@ 4.4 Dynamic Evolution Analysis ($We\in[7,24]$)
$Q=\frac{d_{\textrm{jump}}}{d_{\textrm{max}}},$ (53)
where $d_{\textrm{jump}}$ is the droplet diameter at detachment time $t_{\textrm{contact}}$ and $d_{\textrm{max}}$ represents the global maximum diameter during spreading. As illustrated in Figure 8, the numerical results of pancake quality exhibit excellent agreement with experimental measurements. Figure 8 also shows the temporal characteristics under different Weber numbers, including the detachment time, maximum spreading time and take-off time along with their experimental counterparts. The comparison demonstrates satisfactory consistency between numerical predictions and physical observations.
@@
101015152020252555101015152020WeContact duration (ms)Experimental ($t_{\textrm{contact}}$)SPH ($t_{\textrm{contact}}$)
1010151520202525223344556677WeMoments of maximum spread (ms)Experimental ($t_{\textrm{max}}$)SPH ($t_{\textrm{max}}$)
@@
Figure 8: Comparison of droplet dynamics between SPH simulations and experimental measurements [8]. Error bars represent standard deviations from experimental repetitions.
4.5 Oblique bouncing on tilted substrates ($We=31.2$, $\theta=30^{\circ}$)
@@
4.6 Secondary droplets formed by the Worthington jet ($We=14.1$)
@@
Droplets with moderate Weber numbers have the potential to undergo secondary impacts subsequent to their initial interaction with the substrate. This secondary impact is often more forceful, imparting sufficient momentum to the droplet for it to rebound to a greater height and produce satellite droplets – a phenomenon extensively documented in prior research [50, 51]. To assess the ability of our model to capture subsequent droplet dynamics, we conducted extended numerical simulations for a droplet with We=14.1 (Supplementary Movie 2). As depicted in Figure 10, the numerical outcomes effectively replicate the entire sequence of events: secondary impact, jet formation, and eventual satellite droplet generation. These findings underscore the reliability of the proposed contact algorithm and the precision of the cohesion model integrated in this investigation.
@@ 5 Conclusions
Appendix A Derivation of Equations (12) and (13)
Proof 2
@@ Appendix A Derivation of Equations (12) and (13)
$G_{f}:=\int_{0}^{R_{1}}\int_{0}^{\pi}\int_{0}^{\arccos(s/R_{1})}\int_{s/\cos\theta}^{R_{1}}f(r)\cos\theta\,r^{2}\sin\theta\,drd\theta d\phi ds.$
@@
$G_{f}$ $=\pi\int_{0}^{R_{1}}\int_{0}^{\arccos(s/R_{1})}\cos\theta\sin\theta\int_{s/\cos\theta}^{R_{1}}f(r)r^{2}drd\theta ds.$ (54)
Next, applying the substitution $u=s/R_{1}$ transforms (54) to:
$G_{f}=\pi R_{1}\int_{0}^{1}\int_{0}^{\arccos u}\cos\theta\sin\theta\int_{R_{1}u/\cos\theta}^{R_{1}}f(r)r^{2}drd\theta du.$
@@
$G_{f}=\pi R_{1}\int_{0}^{\pi/2}\cos\theta\sin\theta\int_{0}^{\cos\theta}\int_{R_{1}u/\cos\theta}^{R_{1}}f(r)r^{2}drdud\theta.$
A second exchange of integration order gives:
$G_{f}$ $=\pi R_{1}\int_{0}^{\pi/2}\cos\theta\sin\theta\int_{0}^{R_{1}}f(r)r^{2}\int_{0}^{r\cos\theta/R_{1}}dudrd\theta$
$=\pi\int_{0}^{\pi/2}\cos^{2}\theta\sin\theta d\theta\int_{0}^{R_{1}}f(r)r^{3}dr$
$=\frac{\pi}{3}\int_{0}^{R_{1}}f(r)r^{3}dr,$ (55)
@@
For the special case where $f(|\mathbf{x}|)=\partial_{|\mathbf{x}|}W(\mathbf{x},h)$ with $W$ being a kernel function, we proceed as follows. The normalization condition requires:
$1=\int W(\mathbf{x},R_{1})d\mathbf{x}=4\pi\int_{0}^{R_{1}}\tilde{W}(r)r^{2}dr,$ (56)
where $\tilde{W}(|\mathbf{x}|)=W(\mathbf{x},h)$.
@@
$G_{f}$ $=\frac{\pi}{3}\int_{0}^{R_{1}}\partial_{r}\tilde{W}(r)r^{3}dr$
$=\frac{\pi}{3}\left[\left.r^{3}\tilde{W}(r)\right|_{0}^{R_{1}}-3\int_{0}^{R_{1}}\tilde{W}(r)r^{2}dr\right]$
$=-\pi\int_{0}^{R_{1}}\tilde{W}(r)r^{2}dr.$ (57)

### v2

@@ 4.4 Dynamic Evolution Analysis ($\textit{We}\in[7,24]$)
(51) $Q=\frac{d_{\text{jump}}}{d_{\text{max}}},$
where $d_{\text{jump}}$ is the droplet diameter at detachment time $t_{\text{contact}}$ and $d_{\text{max}}$ represents the global maximum diameter during spreading. As illustrated in Figure 7, the numerical results of pancake quality exhibit excellent agreement with experimental measurements. Figure 7 also shows the temporal characteristics under different Weber numbers, including the detachment time, maximum spreading time and take-off time along with their experimental counterparts. The comparison demonstrates satisfactory consistency between numerical predictions and physical observations.
@@
101015152020252555101015152020WeContact duration (ms)Experimental ($t_{\text{contact}}$)SPH ($t_{\text{contact}}$)
1010151520202525223344556677WeMoments of maximum spread (ms)Experimental ($t_{\text{max}}$)SPH ($t_{\text{max}}$)
@@
Figure 7: Comparison of droplet dynamics between SPH simulations and experimental measurements [liu2014pancake]. Error bars represent standard deviations from experimental repetitions.
4.5 Oblique bouncing on tilted substrates ($\textit{We}=31.2$, $\theta=30^{\circ}$)
@@
4.6 Secondary droplets formed by the Worthington jet ($\textit{We}=14.1$)
@@
Droplets with moderate Weber numbers have the potential to undergo secondary impacts subsequent to their initial interaction with the substrate. This secondary impact is often more forceful, imparting sufficient momentum to the droplet for it to rebound to a greater height and produce satellite droplets – a phenomenon extensively documented in prior research [aria2014physicochemical, zhang2022impact]. To assess the ability of our model to capture subsequent droplet dynamics, we conducted extended numerical simulations for a droplet with We=14.1 (Supplementary Movie 2). As depicted in Figure 9, the numerical outcomes effectively replicate the entire sequence of events: secondary impact, jet formation, and eventual satellite droplet generation. These findings underscore the reliability of the proposed contact algorithm and the precision of the cohesion model integrated in this investigation.
4.7 Sensitivity of the numerical simulation to spatial discretization
To assess the influence of spatial resolution on model performance, we compared the temporal evolution of droplet diameter and height from numerical simulations with experimental data under different particle spacings. It was observed that the droplet radius increases as the initial particle spacing decreases. This trend is attributed to a reduction in numerical artificial viscosity with finer spatial resolution; diminished viscous effects facilitate easier deformation of the droplet. As illustrated in Figure 10(a), the simulated $t_{\max}$ remains approximately 5 ms with increasing spatial resolution. Furthermore, a comparison of the droplet height evolution shows that $t_{\uparrow}$ stays around 3 ms across all tested particle spacings, as seen in Figure 10(b). These results indicate consistent agreement with experimental observations and suggest that the model exhibits low sensitivity to the initial particle spacing.
[…]
Figure 10: Time evolution of droplet diameter and height for $\textit{We}=14.1$ under different mesh resolutions ($\Delta x_{f}$). The results demonstrate the sensitivity of the numerical simulation to spatial discretization.
@@ 5 Conclusions
Appendix A Proof of Lemma 2.1
Proof A.1.
@@
$G_{f}:=\int_{0}^{R_{0}}\int_{0}^{\pi}\int_{0}^{\arccos(s/R_{0})}\int_{s/\cos\theta}^{R_{0}}f(r)\cos\theta\,r^{2}\sin\theta\,drd\theta d\phi ds.$
@@
(52) $G_{f}$ $=\pi\int_{0}^{R_{0}}\int_{0}^{\arccos(s/R_{0})}\cos\theta\sin\theta\int_{s/\cos\theta}^{R_{0}}f(r)r^{2}drd\theta ds.$
Next, applying the substitution $u=s/R_{0}$ transforms (52) to:
$G_{f}=\pi R_{0}\int_{0}^{1}\int_{0}^{\arccos u}\cos\theta\sin\theta\int_{R_{0}u/\cos\theta}^{R_{0}}f(r)r^{2}drd\theta du.$
@@
$G_{f}=\pi R_{0}\int_{0}^{\pi/2}\cos\theta\sin\theta\int_{0}^{\cos\theta}\int_{R_{0}u/\cos\theta}^{R_{0}}f(r)r^{2}drdud\theta.$
A second exchange of integration order yields:
$G_{f}$ $=\pi R_{0}\int_{0}^{\pi/2}\cos\theta\sin\theta\int_{0}^{R_{0}}f(r)r^{2}\int_{0}^{r\cos\theta/R_{0}}dudrd\theta$
$=\pi\int_{0}^{\pi/2}\cos^{2}\theta\sin\theta d\theta\int_{0}^{R_{0}}f(r)r^{3}dr$
(53) $=\frac{\pi}{3}\int_{0}^{R_{0}}f(r)r^{3}dr,$
@@
For the special case where $f(|\boldsymbol{x}|)=\partial_{|\boldsymbol{x}|}W(\boldsymbol{x},h)$ with $W$ being a kernel function, we proceed as follows. The normalization condition requires:
(54) $1=\int W(\boldsymbol{x},R_{0})d\boldsymbol{x}=4\pi\int_{0}^{R_{0}}\tilde{W}(r)r^{2}dr,$
where $\tilde{W}(|\boldsymbol{x}|)=W(\boldsymbol{x},h)$.
@@
$G_{f}$ $=\frac{\pi}{3}\int_{0}^{R_{0}}\partial_{r}\tilde{W}(r)r^{3}dr$
$=\frac{\pi}{3}\left[\left.r^{3}\tilde{W}(r)\right|_{0}^{R_{0}}-3\int_{0}^{R_{0}}\tilde{W}(r)r^{2}dr\right]$
(55) $=-\pi\int_{0}^{R_{0}}\tilde{W}(r)r^{2}dr.$